OpenStudy (lgbasallote):

find the volume generated by rotating the region bounded by \(y^2 - x^2 = 1\) and y = 2 about the x - axis i got \(2\sqrt 3 \pi\) but my book says \(4 \sqrt 3 \pi\) o.O

OpenStudy (lgbasallote):

the equation i used is \[2\pi \int_0^2 y(\sqrt{y^2 - 1}) dy\]

OpenStudy (dumbcow):

using disk or washer method would be easier here cross-sections are rings with outer radius of 2, inner radius of y \[V = \pi \int\limits_{-\sqrt{3}}^{\sqrt{3}}2^{2} -(\sqrt{1+x^{2}})^{2} dx = \pi \int\limits_{-\sqrt{3}}^{\sqrt{3}}(3-x^{2}) dx\]

OpenStudy (dumbcow):

oh i just noticed something, in the limits of your integral....it should be from 1 to 2 not 0 to 2

OpenStudy (lgbasallote):

why 1??

OpenStudy (lgbasallote):

oh wait...it's the graph isnt it

OpenStudy (dumbcow):

because 1 is vertex of hyperbola

OpenStudy (dumbcow):

(0,1) i should say

OpenStudy (lgbasallote):

did that come from the drawing of the graph? i drew it wrong?

OpenStudy (lgbasallote):

oh yeah lol i drew it wrong :p

OpenStudy (lgbasallote):

well the answer to my equation is still \(2\pi\)

OpenStudy (lgbasallote):

i mean \(2\sqrt 3 \pi\)

OpenStudy (dumbcow):

ahh i got it...the height should be 2x or 2sqrt(y^2-1)

OpenStudy (lgbasallote):

but whyyyy T_T

OpenStudy (dumbcow):